OpenStudy (anonymous):

If points A, B, and C lie on a line, they are said to be ______________ points. If the legs of a right triangle are 8 and 10, find the length of the hypotenuse.

OpenStudy (anonymous):

i know the c2 = a2 +B2 64 + 100 = 164 but this doesnt equall a number X itself so not sure what the answer should be

OpenStudy (anonymous):

is it collinear

OpenStudy (anonymous):

okay i got collinear i just need help with the right triangle

OpenStudy (asnaseer):

collinear is correct

OpenStudy (asnaseer):

and: \(\sqrt{164}=\sqrt{4*41}=?\)

OpenStudy (anonymous):

2 square root of 41

OpenStudy (asnaseer):

correct

OpenStudy (anonymous):

thanks for your help

OpenStudy (asnaseer):

yw :)